Identify the conic x^2- 2xy + y^2 + √2x = 2


1
Expert's answer
2021-02-24T07:10:02-0500

Let's solve this problem

first of all we simplify expression

"x^2 - 2 x y + y^2 + \sqrt{2 x} = 2"

"x^2 - 2 x y + \sqrt{2} \sqrt{x} + y^2 = 2"

The graph of conic



Slutions:

Solve for y:

"\sqrt{2}\sqrt{x}+x^2-2xy+y^2 = 2"

we solve quadratic equation by completing the square.

Subtract "\sqrt{2}\sqrt{x}+x^2" from the both sides.

"y^2-2xy = 2-\sqrt{2}\sqrt{x}-x^2"

"D = b^2-4ac = 4x^2+4(2-\sqrt2\sqrt{x}-x^2) = 8 - 4\sqrt{2}\sqrt{x}"

"x_{1,2} = \large\frac{-b\pm \sqrt{D}}{2a}" = "\large\frac{2x\pm\sqrt{8-4\sqrt{2x}}}{2}" "= x \pm \sqrt{2 - \sqrt{2x}}"

Integer solutions: x = 2, y = 2

Implicit derivtives:

"\large\frac{\delta x(y)}{dy} = \frac{4 x - 2 x^3 - 4 y + 6 x^2 y - 6 x y^2 + 2 y^3}{1 + 4 x - 2 x^3 - 4 y + 6 x^2 y - 6 x y^2 + 2 y^3}"

"\large\frac{dy}{\delta x(y)} = \frac{1 + 4 x - 2 x^3 - 4 y + 6 x^2 y - 6 x y^2 + 2 y^3}{4 x - 2 x^3 - 4 y + 6 x^2 y - 6 x y^2 + 2 y^3}"